Please help me solve this:The second of two numbers is 3 less than twice the first.Their sum is 36. Find the numbers. so far i have tried,this:
let x represent the first number
let 3-x(2) represent the second number
x+3-x(2)=36
2x+3(2)=36
2x+6=36
2x=36-6
2x=30
2x divided by 2=30 divided by 2
x=15
15-3(2)=12(2)=24
so for my 2 numbers i have 15 and 24

Answer by checkley71(8403)

CLOSE BUT NO CIGAR:
X+Y=36
NOW DEFINE Y=(2X-3) & REPLACE Y BY (2X-3) & SOLVE FOR THE X VALUE
X+(2X-3)=36
X+2X-3=36
3X=36+3
X=39/3
X=13 ANSWER FOR THE SMALLER NUMBER
2*13-3=26-3=23 FOR THE LARGER NUMBER.
PROOF
13+23=36
36=36 &:
13+(2*13-3)=36
13+26-3=36
13+23=36
36=36